### v1

# Training-time Neuron Alignment through Permutation Subspace for Improving Linear Mode Connectivity and Model Fusion

Abstract
In deep learning, stochastic gradient descent often yields functionally similar yet widely scattered solutions in the weight space even under the same initialization, causing barriers in the Linear Mode Connectivity (LMC) landscape. Overcoming these barriers is crucial for understanding deep learning dynamics and enhancing model-fusion algorithms. Previous studies highlight the role of permutation symmetry in reducing post-training barriers through network permutation. However, these post-hoc methods, demanding extra computations, are less effective for larger, complex models (e.g., ViT, LLM) due to numerous permutation matrices. Thus, in this paper, we study training-time neuron alignment. Our hypothesis suggests that training-time permutation subspace can reduce LMC barriers for free. We find that pruning at initialization supports this. Beyond pruning, we introduce TNA-PFN, a simple yet lossless algorithm using a partial gradient mask during training. TNA-PFN is theoretically and empirically validated for reducing LMC barriers. It excels in wide model fusion applications, especially in federated learning, where two algorithms based on TNA-FPN, named FedPFN and FedPNU, are proposed to show its prospects even under heterogeneous datasets. Moreover, TNA-PFN can enhance the generalization of model soup (Wortsman et al., 2022) for vision transformers and ColD fusion (Don-Yehiya et al., 2023) for pretrained language models.
1 Introduction
Understanding the loss landscape of Deep Neural Networks (DNNs) is the key to understanding the mechanisms and training dynamics behind generalization and optimization (Li et al., 2018b; Fort & Jastrzebski, 2019; Simsek et al., 2021; Vlaar & Frankle, 2022), and it is still an open problem. Empirical findings demonstrate that Stochastic Gradient Descent (SGD) will find many minima that are functionally similar but far away from each other in parameter space (Draxler et al., 2018; Zhang et al., 2021; Entezari et al., 2022). The literature in Linear Mode Connectivity (LMC) suggests that if we linearly interpolate two independently trained networks, which have the same initialization and trainset but have different SGD random seeds (i.e., batch orders), there will be a loss barrier in the landscape (Draxler et al., 2018; Garipov et al., 2018; Ainsworth et al., 2022) (also see Figure 3). The loss barrier reflects that the two SGD solutions fall into two different loss basins that cannot be linearly connected, and it is detrimental to model fusion (Ainsworth et al., 2022; Li et al., 2022).
Figure 1: A simple demonstration of reducing permutation symmetries. There are 6 networks (1 original + 5 counterparts) which are functionally identical but with different permutations. The number of permutation symmetries can be reduced by asymmetrically fixing some weights (in red). Though this demonstration presents a simple static network permutation, we will show in the main paper that this kind of method can realize better neuron alignment under training dynamics.
The reasons behind the barrier in LMC are mainly the over-parameterization and permutation symmetry (also known as permutation invariance) properties of DNNs. Over-parameterization explains why there are abundant minima found by SGD (Zhang et al., 2021; Neyshabur et al., 2017; Safran & Shamir, 2018). Permutation symmetry suggests that the function of the network can remain the same while changing the permutations of neurons, which can result in many functionally same but geometrically different solutions (Ainsworth et al., 2022; Wang et al., 2020a; Tatro et al., 2020). Previous works seek to achieve neuron alignment for improving LMC by utilizing the permutation symmetry property after training. In Entezari et al. (2022), it is conjectured that if taking all permutations into account, all SGD solutions can be mapped into the same loss basin where no barrier in LMC. Git Re-Basin (Ainsworth et al., 2022) further validates this conjecture by proposing three algorithms to find such permutations.
However, post-hoc neuron alignment is a hard combinatorial optimization problem. As stated in (Ainsworth et al., 2022), even for a three-layer MLP with 512 widths, the number of permutation symmetries is nearly 10^3498. When comes to larger foundation models, especially vision transformers (Dosovitskiy et al., 2021) and large language models (Radford et al., 2018), finding such an appropriate permutation to align neurons is more challenging and expensive. For the scenarios where alignment and fusion among multiple models are needed, especially federated learning (Wang et al., 2020b), the cost of the post-hoc alignment methods increases as the number of models.
Can we realize training-time neuron alignment?
If training-time alignment is reached, it will relieve the cost and necessity of post-matching. Also, when the model size and the number of models grow, the training-time method can boost model fusion for free without growing computations. In this paper, we hypothesize the key is to break permutation symmetry via permutation asymmetry subspace. During training, if the potential permutation symmetries are reduced, the neurons will be more aligned in a subspace so that the LMC barriers will decrease and model fusion will improve. A simple demonstration of how permutation asymmetry subspace works is in Figure 1.
We find that pruning at initialization supports our hypothesis. To overcome the limitations of pruning, we introduce TNA-PFN, a simple yet lossless algorithm using a partial gradient mask during training, which is then theoretically and empirically validated for reducing LMC barriers. TNA-PFN excels in wide model fusion applications, including in federated learning under heterogeneous datasets, model soup (Wortsman et al., 2022) for vision transformers, and ColD fusion (Don-Yehiya et al., 2023) for pretrained language models. In a nutshell, we have the following contributions.
  • We discover the neuron alignment problem from the perspective of training time, which provides new insights. We hypothesize that learning in permutation subspaces can reach better LMC.
  • Under the hypothesis, we first find pruning at initialization can improve LMC. Then, we propose a simple yet more lossless training-alignment method TNA-PFN, which is validated under both theoretical and empirical analysis.
  • TNA-PFN excels in wide model fusion applications. We extend TNA-PFN in federated learning and devise two algorithms, FedPFN and FedPNU. Also, TNA-PFN can boost model soup and ColD fusion for free, showcasing its prospects for pretrained foundation models.
2 Background and Related Works
In this section, we provide the basic backgrounds and definitions regarding linear mode connectivity and permutation invariance. Additionally, the preliminary of federated learning is in Appendix D.
Linear mode connectivity. In this paper, we focus on the linear mode connectivity of two SGD solutions, which have the same initialization but different data orders [^1] We note that there are other forms of LMC, such as the LMC from the initialization and the trained model (Vlaar & Frankle, 2022), and the LMC between two models with different initializations (Entezari et al., 2022). While in this paper, we focus on LMC and model fusion with specific applications in federated learning, we only consider the LMC cases with the same initialization.. We present the definitions of loss barrier and accuracy barrier below.
Definition 2.1.
(Loss barrier (Entezari et al., 2022)) Let $f_{\mathbf{w}}(\cdot)$ be a function represented by a neural network with parameter vector $\mathbf{w}$ that includes all parameters and $\mathcal{L}(\mathbf{w})$ be the any given loss (e.g., train or test error) of $f_{\mathbf{w}}(\cdot)$. Given two independently trained networks $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, let $\mathcal{L}(\alpha\mathbf{w}_{1}+(1-\alpha)\mathbf{w}_{2})$, for $\alpha\in[0,1]$ be the loss of the linearly interpolated network. The loss barrier $B_{loss}(\mathbf{w}_{1},\mathbf{w}_{2})$ along the linear path between $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ is defined as the highest difference between the loss of the interpolated network and linear interpolation of the loss values of the two networks:
$B_{loss}(\mathbf{w}_{1},\mathbf{w}_{2})=$ $\sup_{\alpha}\left\{[\mathcal{L}(\alpha\mathbf{w}_{1}+(1-\alpha)\mathbf{w}_{2})]\right.$
$\left.-[\alpha\mathcal{L}(\mathbf{w}_{1})+(1-\alpha)\mathcal{L}(\mathbf{w}_{2})]\right\}.$ (1)
The loss barrier of the above definition is not bounded. To better depict and compare the barrier changes, we then propose a definition of the accuracy barrier which is bounded within $[0,1]$.
Definition 2.2.
(Accuracy barrier) Let $\mathcal{A}(\mathbf{w})$ be the accuracy (e.g., train or test accuracy) of $f_{\mathbf{w}}(\cdot)$. Let $\mathcal{A}(\alpha\mathbf{w}_{1}+(1-\alpha)\mathbf{w}_{2})$, for $\alpha\in[0,1]$ be the accuracy of the linearly interpolated network. The accuracy barrier $B_{acc}(\mathbf{w}_{1},\mathbf{w}_{2})$ along the linear path between $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ is defined as the highest ratio of the interpolated network’s accuracy drop to the averaged accuracy:
$B_{acc}(\mathbf{w}_{1},\mathbf{w}_{2})=\sup_{\alpha}\left[1-\frac{\mathcal{A}(\alpha\mathbf{w}_{1}+(1-\alpha)\mathbf{w}_{2})}{\alpha\mathcal{A}(\mathbf{w}_{1})+(1-\alpha)\mathcal{A}(\mathbf{w}_{2})}\right].$ (2)
The above definition maps the barrier into $[0,1]$. If the accuracy barrier is 0, it means no barrier exists along the linear interpolation path; else if the barrier is nearly 1, it means the generalization of the interpolated model is nearly zero, and its prediction is no better than random guessing.
Permutation invariance. Permutation invariance refers to the property that the positions (i.e., permutations) of neurons of a given network can be changed without changing the network’s function, and it is also known as permutation symmetry (Ainsworth et al., 2022). We take a multi-layer MLP as an example to demonstrate the property.
Assume an MLP network has $L+1$ layers (containing input and output layer), and each layer contains $J_{l}$ neurons, where $l\in\{0,1,\cdots,L\}$ is the layer index. $J_{0}$ and $J_{L}$ are input and output dimensions. We denote the parameters of each layer as the weight matrix $\mathbf{W}_{l}\in\mathbb{R}^{J_{l}\times J_{l-1}}$ and the bias vector $\mathbf{b}_{l}\in\mathbb{R}^{J_{l}}$, $l\in\{1,2,\cdots,L\}$. The input layer does not have parameters. We use $\mathbf{h}_{l}\in\mathbb{R}^{J_{l}}$ as the outputs of the $l$-th layer. We have $\mathbf{h}_{l}=\sigma_{l}(\mathbf{W}_{l}\mathbf{h}_{l-1}+\mathbf{b}_{l})$, where $\sigma_{l}(\cdot)$ is the element-wise activation function, e.g., ReLU. We use $\mathbf{\Pi}\in\{0,1\}^{J\times J}$ as a permutation matrix that satisfies $\sum_{j}\mathbf{\Pi}_{\cdot,j}=1$ and $\sum_{j}\mathbf{\Pi}_{j,\cdot}=1$. By applying the permutation matrices to the layers, the network function remains unchanged. For the $l$-th layer, the layer-wise permutation process is
$\mathbf{h}_{l}=\sigma_{l}(\mathbf{\Pi}_{l}\mathbf{W}_{l}\mathbf{\Pi}_{l-1}^{T}\mathbf{h}_{l-1}+\mathbf{\Pi}_{l}\mathbf{b}_{l}),$ (3)
where $\mathbf{\Pi}_{0}=\mathbf{I}$ and $\mathbf{\Pi}_{L}=\mathbf{I}$, meaning that the input and output are not shuffled. We note that the permutation matrices have the following properties:
$\mathbf{\Pi}^{T}\mathbf{\Pi}=\mathbf{I},\mathbf{\Pi}\mathbf{a}+\mathbf{\Pi}\mathbf{b}=\mathbf{\Pi}(\mathbf{a}+\mathbf{b}),$
$\mathbf{\Pi}\mathbf{a}\odot\mathbf{\Pi}\mathbf{a}=\mathbf{\Pi}(\mathbf{a}\odot\mathbf{b}),\sigma(\mathbf{\Pi}\mathbf{x})=\mathbf{\Pi}\sigma(\mathbf{x}),$ (4)
where $\mathbf{I}$ is the identity matrix, $\odot$ denotes Hadamard product, and $\sigma(\cdot)$ is an element-wise function.
The connection between LMC and permutation invariance. In Entezari et al. (2022); Ainsworth et al. (2022), it is conjectured that if applying appropriate permutation matrices to the networks, two SGD solutions that have barriers before can be mapped into the same loss basin and linearly connected (with low barriers). In previous literature, post-hoc matching methods are proposed to approximate the right permutations (Ainsworth et al., 2022; Peña et al., 2023).
The comparison between our paper and Frankle et al. (2020). In Frankle et al. (2020), the authors find that better LMC will result in better results of the lottery ticket hypothesis (a kind of pruning). However, their conclusions cannot infer that pruning will result in better LMC, where our contributions lie (see more detailed discussions in Appendix E). It is noteworthy that we are the first to present this finding and it is non-trivial.
We include detailed related works in Appendix E.
Figure 2: LMC of random pruning at initialization and TNA-PFN under different mask ratios. For pruning, the mask ratio is the pruning ratio. “Avg. Acc.” means averaged accuracies of individual models, and “Interp. Acc.” means the accuracy of the interpolated model ($\alpha=0.5$) of two modes. The shadow areas mean the accuracy barriers in LMC, the smaller the better.
3 Permutation Subpace and Linear Mode Connectivity
3.1 Hypothesis and Preliminary Finding
Previous works find that permutation symmetry is the main cause of LMC barriers. Due to the numerous parameters and permutation symmetries, SGD will find solutions far from each other in the landscapes during training. Therefore, we think a possible solution for training-time neuron alignment is breaking permutation symmetries in training via the same asymmetric subspace across models, so we make the following hypothesis.
Hypothesis 3.1 (Informal).
If we can reduce the potential number of permutation symmetries by learning different models in the same permutation subspace, the linear mode connectivity will be improved.
Network pruning improves LMC. One straight-forward method which fits Hypothesis 3.1 is network pruning at initialization. We apply random weight pruning to an initialized model, and different neurons will have different pruning structures (i.e., pruning ratios and pruning positions), so the symmetric structure of the network is broken, and the permutations of neurons are limited. From this pruned initialization, different training runs will learn in the same subspace where pruned weights are always zeros and permutations of neurons are constrained.
It is validated in Figure 2 that pruning can actually improve LMC compared with vanilla training, which supports our hypothesis. But when the pruning ratio $\rho$ is high (i.e., 0.8 and 0.9), pruning will result in an untrainable network with nearly zero generalization. It is more severe when the initialized model is pretrained transformers, from Figure 14, it can be seen that even if the pruning ratio is mild ($\rho=0.2$), random pruning is fatal to ViT’s network generalization. Thus, to overcome pruning’s drawbacks, we propose a simple yet lossless (in terms of generalization) training-time alignment method in subsection 3.2.
3.2 Training-time Neuron Alignment with Partially Fixed Neurons
3.2.1 Method Formulation
Instead of setting the weights as zeros like pruning, we propose to fix some neurons’ weights, which will also break the network symmetry to reduce the permutations while preserving the expressiveness power of networks. An intuitive demonstration is in Figure 1, by fixing some weights of neurons, the number of potential permutations decreases.
Specifically, given an initial network parameterized by a weight vector $\mathbf{w}^{0}\in\mathbb{R}^{d}$. For $\mathbf{w}^{0}$, we randomly generate a mask for each layer according to the mask ratio $\rho$ (refers to the proportion of zeros in the mask $\mathbf{m}^{0}$), and the whole mask is $\mathbf{m}^{0}\in\{0,~{}1\}^{d}$. In $\mathbf{m}^{0}$, 00 for fixed and $1$ for updated, indicating the parameter update status. We individually train $n$ models with different batch orders or datasets. We set each model’s initialization as $\mathbf{w}_{i}\leftarrow\mathbf{w}^{0},i\in[n]$. Each model $\mathbf{w}_{i},i\in[n]$ conducts the following updates in every SGD iteration:
$\mathbf{w}_{i}\leftarrow\mathbf{w}_{i}-\eta(\mathbf{m}^{0}\odot\mathbf{g}_{i}(\mathbf{w}_{i})),$ (5)
where $\odot$ denotes the element-wise Hadamard product, $\eta$ refers to the learning rate, and $\mathbf{g}_{i}$ is its gradients of the optimizer, such as SGD or Adam. After training for $E$ epochs, we validate the LMC with respect to the loss or accuracy barriers in Definitions 2.1 and 2.2. The method is notated as TNA-PFN.
We also present how TNA-PFN improves LMC under different $\rho$ in Figure 2, which shows TNA-PFN’s effectiveness in reducing the barriers and its advantages over pruning. Additionally, a connectivity-accuracy tradeoff is observed for both TNA-PFN and pruning that when the mask ratio is higher, the accuracy barriers diminish along with the decrease in the averaged accuracies of independently trained models. However, when $\rho$ is set appropriately (e.g., 0.4-0.6 for the CIFAR-10 and CNN setting), both the averaged accuracy and LMC can be improved.
Discussions on differences between TNA-PFN and previous gradient-mask-based methods. Applying gradient masks is discovered in previous gradient compression literature of distribution optimization, but our method is different from the previous works in the aspects as follows. Motivation difference: Gradient compression is proposed for communication efficiency of distributed optimization while we study the training-time neuron alignment problem in LMC. Implementation difference: Gradient compression uses different random top-k gradient masks at each worker and changes the mask per communication iteration (Lin et al., 2018; Vogels et al., 2019); whereas, TNA-PFN uses the same random gradient masks at each model, fixes the mask, and independently trains the models without any communications; and FedPFN/FedPNU (presented in subsection 4.2) uses the same masks at each client’s local training and changes the mask per global communication round. Effect difference: Since the masks of workers are different and changing, previous gradient compression methods cannot learn in a consistent subspace of parameters, while we learn in a subspace by the same gradient mask so that some neuron weights are not updated.
3.2.2 Theoretical Analysis
We make a theoretical analysis about how TNA-PFN can improve LMC, shown in Theorem 3.2 (proof is in Appendix B). The main idea is to treat the linear interpolated landscape of the barrier as a function of parameter $\alpha$, and the connectivity can be depicted by the first and second derivatives of the function.
Theorem 3.2.
[…]
$\left|z(\alpha)\right|\leq\sqrt{2}b\sigma_{\bm{v}}\sigma_{\bm{U}}\log(8h/\delta)\sqrt{h}\sqrt{1-\rho_{\bm{U}}},$ (6)
$\left|\frac{dz(\alpha)}{d\alpha}\right|\leq 4\sqrt{2}b\sigma_{\bm{v}}\sigma_{\bm{U}}\log{(24h/\delta)}\sqrt{h}(\sqrt{1-\rho_{\bm{v}}}+\sqrt{1-\rho_{\bm{U}}}),$ (7)
$\left|\frac{d^{2}z(\alpha)}{d\alpha^{2}}\right|\leq 8b\sigma_{\bm{v}}\sigma_{\bm{U}}\log(4h/\delta)\sqrt{h}\sqrt{(1-\max\{\rho_{\bm{U}},\rho_{\bm{v}}\})},$ (8)
where $\rho_{\bm{v}}$ and $\rho_{\bm{U}}$ refer to the mask ratios (the proportion of zeros in the mask) of masks ${\bm{M}}_{{\bm{v}}}$ and ${\bm{M}}_{{\bm{U}}}$ respectively.
Remark 3.3.
$\left|z(\alpha)\right|$ is the barrier given $\alpha$. $\left|\frac{dz(\alpha)}{d\alpha}\right|$ demonstrates the barrier function changes along the interpolation path $\alpha\in[0,1]$, and the smaller value means smaller changes. If $\left|\frac{dz(\alpha)}{d\alpha}\right|\to 0$, it means that $z(\alpha)$ is a constant, but it does not mean $z(\alpha)$ is a linear function of $\alpha$. $\left|\frac{d^{2}z(\alpha)}{d\alpha^{2}}\right|$ reflects the linearity of function $z(\alpha)$, and if $\left|\frac{d^{2}z(\alpha)}{d\alpha^{2}}\right|\to 0$, it means that $z(\alpha)$ is linear w.r.t. $\alpha$.
Table 1: The performances of post-matching methods after TNA-PFN. Interpolated Accuracy (Interp. Acc.) means the accuracy of the linearly interpolated model, i.e., $\mathcal{A}(0.5\mathbf{w}_{1}+0.5\mathbf{w}_{2})$. “Iter.” refers to the number of iterations in the post-matching methods, reflecting the computation costs.
CIFAR-10 MNIST
MLP_h2_w200 ResNet20 MLP_h5_w200 MLP_h6_w200
[…]
Interp. Acc. w/o Post-matching 31.9±2.4 43.7±0.4 36.1±4.3 46.2±4.7 59.4±24.2 84.8±8.2 63.7±15.6 87.5±8.9
Interp. Acc. after 10 Iter. of SA 32.2±2.2 43.7±0.4 36.7±3.4 46.2±4.7 59.7±24.2 85.4±8.0 64.9±14.4 87.7±9.1
Interp. Acc. after 100 Iter. of SA 31.9±2.4 43.7±0.4 36.1±4.3 46.2±4.7 60±24.1 86.9±7.6 64.2±15.1 88.2±7.9
Interp. Acc. after WM 44.7±1.3 48.5±0.9 53.7±2.9 53.6±2.5 96.9±0.3 97.1±0.2 96.8±0.3 96.9±0.4
Required Iter. in WM 5.2±1.0 4.8±1.5 4.6±0.5 2.5±0.2 10.4±1.2 7.6±3.8 11.2±1.8 7.33±4.2
Figure 3: Left two: Accuracy barriers of MLP under different hidden layers ($h$) and widths ($w$). Right two: Loss landscapes of MLP. For MLPs, if the barriers exist, TNA-PFN can reduce them. The shadow areas refer to the standard deviations.
Theorem 3.2 shows how TNA-PFN improves LMC by masking some weights from updating (setting $\rho_{\bm{v}}$ and $\rho_{\bm{U}}$ larger than zeros), where the networks learn in a unified permutation subspace.
3.2.3 Empirical Results on Linear Mode Connecitivity
In this subsection, we will conduct experiments to validate the effectiveness of TNA-PFN in improving LMC. If not mentioned otherwise, the mask ratio $\rho$ of TNA-PFN is 0.4 (the hyperparameter which is mild across various settings).
Figure 4: Accuracy barriers under different model architecture. WRN56 abbreviates for WideResNet56. CIFAR-10.
Different model depths, widths, and architectures. In Figure 3, we conduct experiments on MLP with different hidden layers and widths. For MNIST (LeCun & Cortes, 2010), we find shallower and wider networks will not cause barriers, which is consistent with the previous observations (Entezari et al., 2022). For CIFAR-10 (Krizhevsky et al., 2009), the barriers always exist under various depths and widths. Our proposed TNA-PFN can obviously reduce the accuracy barriers from 0.3-0.4 to 0.1, and we also visualize the loss landscapes, which illustrate the barrier reductions. To see more results and illustrations, please refer to subsection C.1.
We study the LMC of simple CNN and ResNets and present the results in Figure 4. ResNets (He et al., 2016) have higher barriers than simple CNN, and the barriers are exacerbated when the networks are deeper or wider. It is suggested that TNA-PFN can lower the barriers under different architectures.
Generally, we observe that TNA-PFN has more dominant advantages when the models are wider, and the observations are: (1) the second figure in Figure 3: for CIFAR-10 when width increases, the barriers of vanilla training go up while the barriers of TNA-PFN go down; (2) Figure 4: the barrier reduction of TNA-PFN is more obvious for WRN56 compared with ResNet56. This observation is consistent with previous literature about permutation symmetry and linear mode connectivity (Entezari et al., 2022).
Table 2: Linear mode connectivity of non-random initializations. Dataset: CIFAR-10. Semi-trained CNN: The initialization is first trained on a disjoint subset of CIFAR-10 for 0.5 epoch. Pretrained ResNet18: Pretrained on ImageNet.
Table 3: Linear mode connectivity of multi-model fusion. The number of models is 5.
Models Metrics Vanilla Train TNA-PFN
Semi-trained CNN Avg. Acc. 64.4±0.7 65.8±0.2
Interp. Acc. 52.8±2.3 63.3±1.2
Acc. Barrier 0.181±0.027 0.0413±0.018
Loss Barrier 0.306±0.058 0.0762±0.038
Pretrained ResNet18 Avg. Acc. 65.3±1.2 69.1±0.5
Interp. Acc. 11.9±0.39 54.9±8.2
Acc. Barrier 0.817±0.003 0.205±0.12
Loss Barrier 0.957±0.45 0.395±0.2
Datasets / Models Metrics Vanilla Train TNA-PFN
CIFAR-10 / CNN Avg. Acc. 63.1±0.6 65.5±0.3
Interp. Acc. 21.3±9.1 48.3±7.2
Acc. Barrier 0.663±0.14 0.264±0.11
CIFAR-10 / MLP_h2_w200 Avg. Acc. 44.2±0.5 48.4±0.5
Interp. Acc. 21.6±1.9 36.6±1.4
Acc. Barrier 0.511±0.043 0.245±0.035
MNIST / MLP_h5_w200 Avg. Acc. 96.5±0.3 96.5±0.2
Interp. Acc. 34.5±15.5 87.1±9.4
Acc. Barrier 0.643±0.16 0.0974±0.096
The role of post-hoc neuron alignment methods after training-time alignment. We consider simulated annealing (SA) (Entezari et al., 2022) and weight matching (WM) (Ainsworth et al., 2022) after TNA-PFN in Table 1. SA requires large computations, and we notice the improvements are also marginal. Under limited computation budgets (10 or 100 iterations), we find that TNA-PFN can reach a higher result than vanilla training after SA. For WM, it is indicated that after TNA-PFN, the required iterations are shortened while the interpolated accuracies are similar. The results reveal that training-time neuron alignment can reduce the costs of post-matching and remain similar or even better post-matched LMC.
Table 4: Results of loss barriers on more deep learning tasks.
[…]
Vanilla Train 0.268±0.061 0.0554±0.047 0.710±0.17
TNA-PFN 0.0381±0.0096 0.0355±0.023 0.375±0.17
More deep learning tasks. We conduct more experiments beyond vision tasks and display the results in Table 4. Polynomial approximation task (Peña et al., 2023; von Oswald et al., 2019): we use an MLP with one hidden layer to approximate the second and third polynomial functions: $y=2x^{2}-1,y=(x-3)^{3}$. Sentiment analysis of text (Liu et al., 2022b): we use an LSTM (Graves & Graves, 2012) to predict the sentiment of IMDb reviews (Maas et al., 2011). It can be seen that the loss barriers are decreased by training-time alignment under both polynomial approximation and sentiment analysis tasks. We also implement the experiments on a large-scale dataset, the subset of ImageNet (Deng et al., 2009; tin, Accessed: 2023). The result is shown in Table 10 of Appendix.
Results under non-random initializations. We also examine whether training-time alignment can help when the initializations are not random, which commonly occurs in the pretraining-finetuning paradigm and federated learning. As presented in subsubsection 3.2.3, TNA-PFN is more advantageous to improve LMC under non-random initializations. TNA-PFN nearly clears the barriers under the setting of CNN and CIFAR-10. For the pretrained ResNet18, we observe that TNA-PFN reduces the barrier by a large margin while improving individual accuracies.
Layer-wise analysis. We conduct a layer-wise analysis of TNA-PFN to see which layer matters most in improving LMC in Figure 5 of Appendix.
Table 3: Linear mode connectivity of multi-model fusion. The number of models is 5.